(a–4)(a–2) = a^2 –8 ????????????????????????????????????

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

LHS = (a–4)(a–2) = a(a–2)–4(a–2) = a^2–2a–4a+8
= a^2–6a+8
≠a^2-8 = RHS
The correct statement is (a–4)(a–2) = a^2–6a+8

Thanks